A politician claims that he will get at least 70% of the votes. Out of 300 randomly sampled registered voters, 200 said they will vote for the said politician. Test the claim using 0.10 level of significance.


1
Expert's answer
2022-06-02T12:41:59-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\ge0.7"

"H_a:p<0.7"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Evidence:

Based on the information provided, the significance level is "\alpha = 0.10

," and the critical value for a left-tailed test is "z_c = -1.2816."

The rejection region for this left-tailed test is "R = \{z: z < -1.2816\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{200/300-0.7}{\sqrt{\dfrac{0.7(1-0.7)}{300}}}=-1.26"

Since it is observed that "z = -1.26>-1.2816= z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(Z<-1.26)= 0.103835," and since "p=0.103835>0.10=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.7, at the "\alpha = 0.10" significance level.